# Anomalous Energy Injection in the Gross-Pitaevskii Framework for Turbulence in Neutron Star Glitches

### v2

@@ Abstract
Modeling the depinning avalanche mechanism poses significant challenges. Vortex depinning initiates when a vortex, initially trapped by a pinning site, overcomes the attractive force and begins moving in the direction of the applied superfluid flow. Several factors, such as the critical velocity of the superfluid flow and the size and shape of the pinning site, affect the overall dynamics of depinning [10]. However, an alternative analysis suggests that glitches emerge due to a quantum transition in which the core transforms into a superconducting gluon-quark superfluid state, with resultant vortex depinning occurring at the boundary layer between the core and the surrounding dissipative medium [3]. In this Letter, we demonstrate anomalous energy transfer from quantum pressure to incompressible kinetic energy, thereby proposing a candidate mechanism for turbulent dynamics within a Gross-Pitaevskii analog of neutron star glitches.
@@
The Model: We model the interior of a spinning-down neutron star’s superfluid core using a quasi-two-dimensional (quasi-2D) atomic Bose-Einstein condensate. This rotating BEC system, incorporating dissipative interactions, is described by the damped Gross-Pitaevskii equation (dGPE), presented in its non-dimensional form in Ref. [33].
@@
The time-dependent rotation frequency $\Omega(t)$ describes the rotational deceleration profile of the condensate as follows:
$\Omega(t)=\begin{cases}\Omega_{0}\cos^{2}\left(\frac{\pi t}{2t_{s}}\right),&\text{if }t\leq t_{s},\\ 0,&\text{if }t>t_{s},\end{cases}$ (2)
where $\Omega_{0}$ is the initial rotation frequency of the condensate, and $t_{s}$ is the spin-down time. The $z$-component of the angular momentum is given by $L_{z}=\mathrm{i}\hbar(y\partial_{x}-x\partial_{y})$. In general, glitch events arise from transitions between two distinct non-zero rotational frequencies [35, 16]. However, a partial spin-down can retain vortices (see supplementary Fig.2), which may suppress the amplitude of glitches. While introducing a sudden deceleration can overcome this suppression, it often results in an unstable and unconfined condensate. Therefore, to enhance the vortex avalanche process without destabilizing the condensate, we have instead chosen to fully spin down the condensate.
@@
where $\tilde{V}(x_{\theta},y_{\theta})=3+\cos(n_{\mathrm{crust}}x_{\theta})+\cos(n_{\mathrm{crust}}y_{\theta})$, and the rotated coordinates are $x_{\theta}=\cos(\theta(t))x_{p}+\sin(\theta(t))y_{p}$ and $y_{\theta}=-\sin(\theta(t))x_{p}+\cos(\theta(t))y_{p}$. Here, $n_{\mathrm{crust}}$ determines the number of pinning sites, $r_{\mathrm{crust}}$ is the radius at which $V_{\mathrm{crust}}$ is maximized, and $\Delta r_{\mathrm{crust}}$ represents the crust thickness. The crust potential parameters are based on Refs. [15, 14]. The centrifugal potential, $V_{\mathrm{cent}}(\mathbf{r},t)=\frac{1}{2}\Omega(t)^{2}r^{2}$, ensures uniform condensate density under rotation. Following Refs. [33, 36, 37], we adopt a coherent length scale (comparable to vortex core size) of $\xi=10~\mathrm{fm}$, yielding an energy scale of $\epsilon=207~\mathrm{keV}$. With these parameters, the time scale is calculated as $\tau=3.2\times 10^{-21}~\mathrm{s}$. We compute the kinetic energy components and their spectra based on the analysis of their numerical implementation in Ref. [38]. It is important to note that while the GPE model is inherently limited to describing weakly interacting BECs in relatively small systems, resulting in a mismatch with the microscopic length scales present in neutron stars, it effectively captures vortex depinning phenomena, which are considered strong candidates for characterizing glitch events. Although scaling the vortex population, condensate radius and rotation frequency to astrophysical scales is unfeasible, the GPE model has reproduced Self-Organized Criticality of the pulsar glitch behavior  [16, 15] The simulation involves numerically solving the GPE using the split-step Crank-Nicolson method [39, 40] in a computational domain of size $512\times 512$, with spatial steps $dx=dy=0.05$ and a time step $dt=10^{-3}$ to ensure numerical stability and convergence. The numerical simulations are performed on an NVIDIA A100 GPU using CUDA C codes developed based on [41]. Initially, we prepare the condensate and crust with an initial rotation frequency $\Omega_{0}=2$ through imaginary-time iterations ($t\to-\mathrm{i}t$). This converged profile is then spun down at varying rates, characterized by the spin-down time $t_{s}$, in real-time. We first consider a case with a static crust and no dissipation ($\gamma=0$) while exploring the effects of damping in subsequent analyses.
@@
Further, the vortex avalanche triggered by Magnus flow results in an instantaneous recoupling of the superfluid and normal components, accompanied by a rapid transfer of angular momentum. This abrupt transition in superfluid systems may induce turbulence but cannot be confirmed using density profiles alone. To characterize the turbulent flow in the system, we therefore analyze the incompressible kinetic energy spectra of the condensate.
@@
As shown in Fig. 3, the vortex density exhibits a distinct scaling behavior with respect to time, further confirming the turbulent behavior of the condensate. During the initial Kolmogorov turbulent regime, the vortex-line density $l_{v}$ scales as $t^{-3/2}$, transitioning to $t^{-1}$ scaling in the later Vinen turbulent regime. This temporal decay of vortex-line density is attributed to vortex breakdown in the Kolmogorov regime, as well as to sound radiation (Kelvin-wave emission) and vortex reconnection in the Vinen regime [42, 43, 44]. For shorter spin-down times, the $t^{-3/2}$ scaling persists over a longer timescale, corresponding to a broader spatial extent of the $k^{-3/2}$ scaling in the incompressible kinetic energy spectrum (Fig. 2). Due to the intrinsic scales of the Gross–Pitaevskii framework, the power-law behavior observed here is limited to laboratory-scale systems. However, the functional form of this scaling, the $k^{-5/3}$ spectrum, is consistent with the incompressible turbulence spectrum reported in a distinct astrophysical context [25, 26]. This similarity in spectral behavior, despite the vast difference in physical scales and systems, suggests a common underlying turbulent mechanism may be at play.
@@
The damping coefficient $\gamma$ is used in atomic BECs to model vortex-sound interactions [46]. A comparable approach is employed in neutron star interiors to describe interactions with the outer crust and proton superconducting core. In addition, the impact of $\gamma$ on vortex and turbulent dynamics under the dGPE model provides an important clue to the contribution of the interaction between neutron star interiors, which are mainly composed of superfluid and proton superconductors. For atomic BECs, the damping parameter $\gamma$ can be controlled by varying the temperature of the condensate, significantly affecting the vortex and density-wave dynamics. The damping parameter $\gamma$ is controlled by varying the temperature of the condensate, significantly affecting vortex and density-wave dynamics.
Figure 5: Variation of (a) incompressible kinetic energy spectra and (b) compressible kinetic energy spectra for different values of the damping coefficient $\gamma$. Both spectra are calculated for the case $t_{s}=10$ and averaged over the time range $t=10$ to $t=30$. The spectra exhibit a better fit to turbulent scaling for smaller $\gamma$ values.
@@
Figure 6: Time evolution of the exchange energy between the incompressible and quantum pressure components for spin-down time $t_{s}=10$ (dark-gray vertical line). The different curves represent the behavior of the profile under varying damping strengths $\gamma$. Notably, the injection of incompressible energy is maximized for an optimal damping coefficient, corresponding to neither underdamped nor overdamped conditions.
@@
Figure 5 illustrates the evolution of incompressible and compressible kinetic energy spectra for various values of the damping coefficient $\gamma$. For the strongly turbulent case with $t_{s}=10$ [see Figs. 2(a)], the spectra exhibit pronounced $k^{-5/3}$ scaling in the incompressible component and $k$ scaling in the compressible component for smaller damping values (underdamped regime). The $k$ scaling in the infrared region and $k^{-7/2}$ scaling in the ultraviolet regime for the compressible spectra indicates the presence of weak wave turbulence [21, 24, 26]. At approximately $\gamma\approx 10^{-2}$, the spectra abruptly lose their scaling behavior as stronger damping (overdamped regime) dissipates the turbulent flow. This turbulent flow is further confirmed by the emergence of a prominent $k^{-3}$ scaling and the diminishing $k^{-5/3}$ scaling in the overdamped case ($\gamma\approx 10^{-2}$), indicating the inhibition of vortex breakdown and the formation of a stable vortex lattice structure. It is worthwhile to mention mainly an over-damped condensate ($\gamma>0.02$) has been considered to model superfluid cores of neutron stars due to several potential dissipation pathways [13, 9].
@@
Similar to the non-damped case, we complement the spectral analysis with the exchange energy profile in the presence of damping Fig. 6. The injection of incompressible kinetic energy reaches a maximum at an optimal value of $\gamma$, situated between the underdamped and overdamped regimes. Under these optimal damping conditions, the transfer of energy from quantum pressure to incompressible modes appears to be significantly enhanced. Given that $\gamma$ serves as a simplified parametrization of dissipation in quantum fluids, the underlying mechanism responsible for the observed maximum in incompressible energy injection at this intermediate damping requires further investigation. In the overdamped regime, around $\gamma\approx 10^{-2}$, the positive peak in exchange energy is absent, consistent with the suppression of turbulence under such strongly dissipative conditions.
@@
Summary and Conclusion: We investigated two-dimensional quantum turbulence in a spinning-down condensate subjected to a crust potential that provides vortex pinning sites. By analyzing a continuous spin-down profile, we demonstrated that as the spin-down duration $t_{s}$ approaches the quasi-discrete limit, the condensate exhibits turbulent behavior. Following the spin-down duration $t_{s}$, the neutron star core displays turbulence characterized by an incompressible kinetic energy spectrum with a $k^{-5/3}$ scaling, indicative of the Kolmogorov regime, and a compressible energy spectrum with a $k^{1}$ scaling, consistent with thermalization. The turbulent core subsequently transitions to the Vinen regime, marked by a $k^{-1}$ scaling, where isolated vortices dominate the flow. In both regimes, the scaling behavior weakens as the spin-down duration increases. The vortex-line density $l_{v}$ exhibits a power-law decay, scaling as $t^{-3/2}$ in the Kolmogorov regime and as $t^{-1}$ in the Vinen regime.
@@
The interplay between kinetic energy components reveals that spinning down induces an energy transfer from the incompressible component to quantum pressure. After the spin-down, a reverse transfer from quantum pressure to incompressible energy occurs, which is sufficient to trigger a turbulent cascade after the cessation of rotational forcing. This injection of kinetic energy is indirectly observed in the vortex-line decay profile as an increase in vortex-line density around $t_{s}$, resulting from the transfer of energy back into the incompressible component. This behavior is attributed to the depinning of vortex lattice, leading to a vortex avalanche [47, 9, 48].
@@
The variation of the Kolmogorov spectrum and compressible spectrum with the damping parameter $\gamma$ shows a sharp transition from non-turbulent to turbulent scaling behavior occurs near $\gamma\sim 10^{-2}$. We have also observed an optimal damping coefficient, where the transfer from quantum pressure to incompressible energy is maximized. Under longer spin-down durations, Liu et al. [33] have reported pulsar glitches triggered by vortex avalanches. As quantum turbulence has been shown to suppress collective excitations in condensates [27, 28], the parameters determining the onset of turbulence presented here may be relevant to glitch formation in the presence of a crust potential. While vortex depinning is commonly considered the mechanism behind glitches in crust-bound systems, self-gravitating BECs without pinning potentials can also exhibit vortex expulsion to the periphery when the total circulation exceeds a critical threshold $\xi_{c}$ [49, 26]. Such expulsions could similarly interact with an external crust and trigger glitch-like behavior in the absence of depinning.
The onset of turbulence in BECs with pinning potentials during spin-down provides crucial insight into the prevalent GPE models of neutron star cores [15, 14, 33]. The dependence of turbulence dynamics on spin-down profiles and dissipative interactions within our Gross-Pitaevskii analog allows us to map the model’s parameter space. This mapping is crucial for assessing the theoretical feasibility and inherent limitations of the analogy itself, thereby clarifying which aspects of neutron star glitch phenomenology it may, or may not, be suited to explore [33]. Furthermore, this study is valuable for exploring inhomogeneous two-dimensional quantum turbulence and its associated dynamics, particularly in relation to the emergence of secondary energy injection mechanisms. It is worth noting that in systems with a large number of vortices, secondary energy injection is primarily governed by the transfer of energy from compressible to incompressible modes. This process arises due to vortices traveling longer distances, repeatedly pinning and depinning at various locations. Similar to the partial spin-down scenario, this results in a substantial retention of vortices within the condensate, which inhibits the transfer of energy from quantum pressure to incompressible kinetic modes.
In conclusion, although the Gross-Pitaevskii analog necessarily simplifies the fermionic nature of neutron Cooper-pair superfluidity in neutron-star interiors—including density-dependent 1S0 and 3P2 pairing, possible hyperon contributions, and uncertainties in nuclear many-body physics—it remains the most microscopically faithful effective theory for studying the dynamics of individual quantized vortices and their mutual interactions. While directly applicable only to finite, laboratory-scale systems, it provides essential insight into the fundamental quantum mechanisms that may underpin vortex-mediated phenomena in astrophysical superfluids [33]. These additional physical contributions could, in principle, be incorporated through multi-component or binary-fluid extensions of the GP framework, which offers a promising direction for future research.
Our simulations reveal a previously unreported transient anomalous energy-injection mechanism, driven by quantum pressure, that operates during the highly turbulent phase immediately following major vortex avalanches. Although quantitative extrapolation to kilometre-scale neutron stars is not yet possible, and the effect diminishes under prolonged differential rotation, the consistent appearance of avalanche-like events, power-law glitch-size distributions, and sub-second rise times across numerous independent Gross-Pitaevskii studies supports the astrophysical relevance of the underlying vortex dynamics.
The anomalous energy-injection mechanism identified in the Gross–Pitaevskii analog system, therefore, represents a new and potentially relevant ingredient for future multi-scale and hybrid models of neutron star interiors.

### v3

@@ Abstract
Modeling the depinning avalanche mechanism poses significant challenges. Vortex depinning initiates when a vortex, initially trapped by a pinning site, overcomes the attractive force and begins moving in the direction of the applied superfluid flow. Several factors, such as the critical velocity of the superfluid flow and the size and shape of the pinning site, affect the overall dynamics of depinning [10]. However, an alternative analysis suggests that glitches emerge due to a quantum transition in which the core transforms into a superconducting gluon-quark superfluid state, with resultant vortex depinning occurring at the boundary layer between the core and the surrounding dissipative medium [3]. In this Letter, we demonstrate anomalous energy transfer from quantum pressure to incompressible kinetic energy, thereby proposing a candidate mechanism for turbulent dynamics within a Gross-Pitaevskii (GP) analog of neutron star glitches.
@@
The Model: We model the interior of a spinning-down neutron star’s superfluid core using a quasi-two-dimensional (quasi-2D) atomic Bose-Einstein condensate. This rotating BEC system, incorporating dissipative interactions, is described by the damped Gross-Pitaevskii equation, presented in its non-dimensional form in Ref. [33].
@@
The time-dependent rotation frequency $\Omega(t)$ denotes the rotational deceleration profile of the condensate which have the forms as $\Omega(t)=\Omega_{0}\cos^{2}\left(\frac{\pi t}{2t_{s}}\right)$ for $t\leq t_{s}$ and zero for $t>t_{s}$. Here $\Omega_{0}$ is the initial rotation frequency of the condensate, and $t_{s}$ is the spin-down time. The $z$-component of the angular momentum is given by $L_{z}=\mathrm{i}\hbar(y\partial_{x}-x\partial_{y})$. In general, glitch events arise from transitions between two distinct non-zero rotational frequencies [35, 16]. However, a partial spin-down can retain vortices (see Fig. 6 in Appendix A), which may suppress the amplitude of glitches [36]. While introducing a sudden deceleration can overcome this suppression, it often results in an unstable and unconfined condensate. Therefore, to enhance the vortex avalanche process without destabilizing the condensate, we have instead chosen to fully spin down the condensate.
@@
where $\tilde{V}(x_{\theta},y_{\theta})=3+\cos(n_{\mathrm{crust}}x_{\theta})+\cos(n_{\mathrm{crust}}y_{\theta})$, and the rotated coordinates are $x_{\theta}=\cos(\theta(t))x_{p}+\sin(\theta(t))y_{p}$ and $y_{\theta}=-\sin(\theta(t))x_{p}+\cos(\theta(t))y_{p}$. Here, $n_{\mathrm{crust}}$ determines the number of pinning sites, $r_{\mathrm{crust}}$ is the radius at which $V_{\mathrm{crust}}$ is maximized, and $\Delta r_{\mathrm{crust}}$ represents the crust thickness. The crust potential parameters are based on Refs. [15, 14]. The centrifugal potential, $V_{\mathrm{cent}}(\mathbf{r},t)=\frac{1}{2}\Omega(t)^{2}r^{2}$, ensures uniform condensate density under rotation. Following Refs. [33, 37, 38], we adopt a coherent length scale (comparable to vortex core size) of $\xi=10~\mathrm{fm}$, yielding an energy scale of $\epsilon=207~\mathrm{keV}$. With these parameters, the time scale is calculated as $\tau=3.2\times 10^{-21}~\mathrm{s}$. We compute the kinetic energy components and their spectra based on the analysis of their numerical implementation in Ref. [39]. It is important to note that while the GP model is inherently limited to describing weakly interacting BECs in relatively small systems, resulting in a mismatch with the microscopic length scales present in neutron stars, it effectively captures vortex depinning phenomena, which are considered strong candidates for characterizing glitch events. Although scaling the vortex population, condensate radius and rotation frequency to astrophysical scales is unfeasible, the GP model has reproduced Self-Organized Criticality of the pulsar glitch behavior  [16, 15] The simulation involves numerically solving the GP equation (1) using the split-step Crank-Nicolson method [40, 41] in a computational domain of size $512\times 512$, with spatial steps $dx=dy=0.05$ and a time step $dt=10^{-3}$ to ensure numerical stability and convergence. The numerical simulations are performed on an NVIDIA A100 GPU using CUDA C codes developed based on [42]. Initially, we prepare the condensate and crust with an initial rotation frequency $\Omega_{0}=2$ through imaginary-time iterations ($t\to-\mathrm{i}t$). This converged profile is then spun down at varying rates, characterized by the spin-down time $t_{s}$, in real-time. We first consider a case with a static crust and no dissipation ($\gamma=0$) while exploring the effects of damping in subsequent analyses.
@@
Further, the vortex avalanche triggered by Magnus flow results in an instantaneous recoupling of the superfluid and normal components, accompanied by a rapid transfer of angular momentum. However, for system with large number of vortices the expulsion happens slowly [36]. This abrupt transition in superfluid systems may induce turbulence but cannot be confirmed using density profiles alone. To characterize the turbulent flow in the system, we therefore analyze the incompressible kinetic energy spectra of the condensate.
@@
As shown in Fig. 3, the vortex density exhibits a distinct scaling behavior with respect to time, further confirming the turbulent behavior of the condensate. During the initial Kolmogorov turbulent regime, the vortex-line density $l_{v}$ scales as $t^{-3/2}$, transitioning to $t^{-1}$ scaling in the later Vinen turbulent regime. This temporal decay of vortex-line density is attributed to vortex breakdown in the Kolmogorov regime, as well as to sound radiation (Kelvin-wave emission) and vortex reconnection in the Vinen regime [43, 44, 45]. For shorter spin-down times, the $t^{-3/2}$ scaling persists over a longer timescale, corresponding to a broader spatial extent of the $k^{-3/2}$ scaling in the incompressible kinetic energy spectrum (Fig. 2). Due to the intrinsic scales of the Gross–Pitaevskii framework, the power-law behavior observed here is limited to laboratory-scale systems. However, the functional form of this scaling, the $k^{-5/3}$ spectrum, is consistent with the incompressible turbulence spectrum reported in a distinct astrophysical context [25, 26] and system with large number of vortices [13, 36]. This similarity in spectral behavior, despite the vast difference in physical scales and systems, suggests a common underlying turbulent mechanism may be at play.
@@
The damping coefficient $\gamma$ is used in atomic BECs to model vortex-sound interactions [47]. A comparable approach is employed in neutron star interiors to describe interactions with the outer crust and proton superconducting core. In addition, the impact of $\gamma$ on vortex and turbulent dynamics under the damped GP model provides an important clue to the contribution of the interaction between neutron star interiors, which are mainly composed of superfluid and proton superconductors. For atomic BECs, the damping parameter $\gamma$ can be controlled by varying the temperature of the condensate, significantly affecting the vortex and density-wave dynamics. The damping parameter $\gamma$ is controlled by varying the temperature of the condensate, significantly affecting vortex and density-wave dynamics.
@@
Figure 5: Variation of (a) incompressible kinetic energy spectra, (b) compressible kinetic energy spectra, and (c) Time evolution of the exchange energy between the incompressible and quantum pressure components for spin-down time $t_{s}=10$ (dark-gray vertical line), for different values of the damping coefficient $\gamma$.
@@
Figure 5(a-b) illustrates the evolution of incompressible and compressible kinetic energy spectra for various values of the damping coefficient $\gamma$. For the strongly turbulent case with $t_{s}=10$ [see Figs. 2(a)], the spectra exhibit pronounced $k^{-5/3}$ scaling in the incompressible component and $k$ scaling in the compressible component for smaller damping values (underdamped regime). The $k$ scaling in the infrared region and $k^{-7/2}$ scaling in the ultraviolet regime for the compressible spectra indicates the presence of weak wave turbulence [21, 24, 26]. At approximately $\gamma\approx 10^{-2}$, the spectra abruptly lose their scaling behavior as stronger damping (overdamped regime) dissipates the turbulent flow. This turbulent flow is further confirmed by the emergence of a prominent $k^{-3}$ scaling and the diminishing $k^{-5/3}$ scaling in the overdamped case ($\gamma\approx 10^{-2}$), indicating the inhibition of vortex breakdown and the formation of a stable vortex lattice structure. It is worthwhile to mention mainly an over-damped condensate ($\gamma>0.01$) has been considered to model superfluid cores of neutron stars due to several potential dissipation pathways [13, 9].
@@
Similar to the non-damped case, we complement the spectral analysis with the exchange energy profile in the presence of damping Fig. 5(c). The injection of incompressible kinetic energy reaches a maximum at an optimal value of $\gamma$, situated between the underdamped and overdamped regimes. Under these optimal damping conditions, the transfer of energy from quantum pressure to incompressible modes appears to be significantly enhanced. Given that $\gamma$ serves as a simplified parametrization of dissipation in quantum fluids, the underlying mechanism responsible for the observed maximum in incompressible energy injection at this intermediate damping requires further investigation. In the overdamped regime, around $\gamma\approx 10^{-2}$, the positive peak in exchange energy is absent, consistent with the suppression of turbulence under such strongly dissipative conditions.
@@
Summary and Conclusion: We investigated two-dimensional quantum turbulence in a spinning-down condensate subjected to a crust potential that provides vortex pinning. Continuous spin-down profiles show that as the spin-down duration $t_{s}$ approaches the quasi-discrete limit, turbulent behavior emerges. After spin-down, the system exhibits an incompressible kinetic energy spectrum with Kolmogorov scaling $k^{-5/3}$ and a compressible spectrum scaling as $k^{1}$, indicating thermalization. The flow subsequently enters the Vinen regime, dominated by isolated vortices and characterized by a $k^{-1}$ scaling. In both regimes, spectral scaling weakens as $t_{s}$ increases, while the vortex-line density $l_{v}$ decays as $t^{-3/2}$ (Kolmogorov) and $t^{-1}$ (Vinen).
@@
The interplay between kinetic energy components reveals that spin-down drives a transfer from incompressible kinetic energy to quantum pressure. After spin-down, a reverse transfer from quantum pressure back to incompressible energy triggers a turbulent cascade even in the absence of external forcing. This secondary energy injection appears as an increase in vortex-line density near $t_{s}$ and is attributed to vortex depinning and avalanche events in the neutron star cores [48, 9, 49].
@@
A sharp transition from non-turbulent to turbulent scaling occurs near a damping parameter $\gamma\sim 10^{-2}$, with an optimal value maximizing energy transfer from quantum pressure to incompressible modes. For longer spin-down durations, vortex avalanches have been associated with pulsar glitches [33]. Since quantum turbulence suppresses collective excitations [27, 28], the turbulence onset conditions identified here may be relevant to glitch formation in pinned systems. In addition to depinning-driven glitches, self-gravitating BECs without pinning can expel vortices once the circulation exceeds a critical threshold $\xi_{c}$, potentially producing glitch-like behavior via crust interaction [50, 26].
These findings complement the already existing Gross-Pitaevskii models of neutron star cores [15, 14, 33]. Mapping turbulence dynamics across spin-down profiles and dissipation parameters highlights both the applicability and limitations of the GP analogy for neutron star glitch phenomenology [33]. The results also advance understanding of inhomogeneous two-dimensional quantum turbulence, highlighting secondary energy injection driven by repeated vortex pinning and depinning.
Although the GP framework simplifies neutron superfluid microphysics, it remains the most microscopically faithful model for quantized vortex dynamics. While extrapolation to neutron star scales is not yet quantitative, the robust appearance of vortex avalanches, power-law glitch statistics, and rapid rise times across GP studies supports the astrophysical relevance of the underlying mechanisms. The transient quantum-pressure-driven energy-injection process identified here thus represents a potentially important ingredient for future multi-scale models of neutron star interiors.
Appendix A Energy Injection in Larger Condensates with More Vortices
In this appendix, we provide additional details on the energy injection mechanism in larger condensates supporting a greater number of vortices. To further investigate this mechanism in larger condensates containing a greater number of vortices (e.g., approximately $70$), we performed simulations involving a partial spin-down of the condensate, reducing the frequency from $\Omega=2.0$ to $\Omega=1.0$. Under these conditions, we observed that vortex expulsion does not occur simultaneously; instead, vortices are shed progressively.
Figure 6: Snapshots of condensate density at various times for a larger condensate rotating from $\Omega=2.0$ to $\Omega=1.0$ with spin-down time $t_{s}=10$.
Figure 6 presents snapshots of the condensate density at various time instances during a spin-down of duration $10$ units. In this process, vortices near the condensate center must travel longer distances to reach the edge, increasing the likelihood of multiple pinning–depinning events. This behavior significantly inhibits the vortex avalanche mechanism, thereby reducing the size of glitches in condensates with larger vortex populations.
Figure 7: Vortex fraction of the condensate after times $t=100,75$ and $t=50$, with respect to (a) initial vortex population for a condensate and (b) final rotation frequency $\Omega_{\mathrm{f}}$. In both cases the vortex fraction is higher at later duration, indicating a inhibition of vortex expulsion.
This inhibition is further corroborated by the ratio of vortices at time $t$ ($N_{v_{t}}$)to the initial vortex population ($N_{v_{i}}$), whose evolution with respect to the initial vortex number is illustrated in Fig. 7(a). Since Bose-Einstein condensates (BECs) maintain a uniform areal density, the vortex population serves as a reliable proxy for condensate size. Figure 7(a) shows a linear increase in vortex retention with increasing condensate size. This observation provides a physical explanation for the upper bound on the number of vortices required to observe glitch-like behavior, as reported in Ref. [12].
Additionally, the figure indicates that the vortex fraction at $t=75$ and $t=100$ is nearly identical, suggesting that the vortex expulsion process significantly slows at later times. In Fig. 7(b), the vortex fraction for a larger condensate exhibits an exponential increase with respect to the final rotation frequency $\Omega_{\mathrm{f}}$. For smaller $\Omega_{\mathrm{f}}$, the vortex fraction remains nearly constant, indicating that vortex expulsion is comparably effective for both complete spin-down ($\Omega_{\mathrm{f}}$ = 0) and partial spin-down to lower rotation frequencies ($\Omega_{\mathrm{f}}<0.5$ ).
Despite the inhibited avalanche activity, the larger vortex system still exhibits turbulence before reaching the final rotation frequency. Figure 8(a) shows the incompressible kinetic energy spectra, averaged over the interval $t=5$ to $t=10$, displaying $k^{-5/3}$ and $k^{-3}$ scaling, characteristic of the Kolmogorov cascade.
Figure 8: (a) Incompressible kinetic energy spectra displaying $k^{-5/3}$ and $k^{-3}$ scaling averaged over time $t=5$ to $t=10$, indicating Kolmogorov cascade in the larger condensate with partial spin-down. (b)Vortex density (per unit area) as a function of time displaying a scaling behavior as $t^{-3/2}$ consistent with the Kolmogorov-like cascade.
Figure 8(b) shows the time evolution of vortex density on a logarithmic scale. Similar to the smaller condensate case, a $t^{-3/2}$ scaling is observed during the turbulent regime. However, after this phase, vortex density decays more slowly than in smaller systems, indicating a prolonged persistence of vortices following spin-down. This slow decay further supports the conclusion that inter-vortex spacing and vortex core size do not directly influence the turbulent dynamics in our Gross–Pitaevskii model, as the vortex areal density remains of the same order of magnitude across different system sizes.
It is important to note that turbulence arises before the completion of the spin-down.
Figure 9: Evolution of energy exchange between compressible-incompressible components (red dashed line), quantum pressure-incompressible components (green dash-dotted line), and quantum pressure-compressible components (blue circled line). For larger condensates, transitioning between two non-zero frequencies, compressible to incompressible exchange dominates over the quantum pressure to incompressible exchange due to the persistence of vortices.
This is attributed to the larger number of vortices and the enhanced compressible flow in the system, which together increase the likelihood of de-pinning events occurring before the target rotation frequency is reached—unlike in the smaller condensate case.
To further analyze the energy dynamics during this process, we plot the evolution of energy exchange between the compressible and incompressible components, the quantum pressure and incompressible components, and the quantum pressure and compressible components in Fig. 9. For larger condensates undergoing a transition between two non-zero rotation frequencies, energy transfer from the compressible to the incompressible component dominates over the quantum pressure–to–incompressible transfer. Notably, this energy injection from the compressible flow coincides with the onset of the Kolmogorov cascade, indicating its key role in driving turbulence. The suppression of quantum pressure–to–incompressible transfer in this regime is primarily due to the persistence of vortices and the enhanced generation of sound waves, both of which are facilitated by the larger condensate size.
In contrast, smaller condensates generate fewer sound waves, and vortex expulsion is more pronounced during the spin-down process. As a result, the compressible–to–incompressible transfer is weaker or absent during the turbulent phase, making the quantum pressure the dominant source of incompressible energy injection. Furthermore, vortex expulsion in larger condensates is significantly suppressed, even when the condensate is fully spun down, due to the system’s extended spatial scale. Vortices persist for long durations and are only expelled gradually over time. Since Bose-Einstein condensates sustain vortices only above a certain threshold rotation frequency, a rapidly rotating condensate can remain vortex-free if the interaction strength and trapping potential are tuned appropriately. In such cases, the system can effectively behave like a non-rotating condensate.
Our simulations provide clear evidence of the simultaneous onset of turbulence and energy injection before the system reaches the lower rotation frequency. Most importantly, incompressible energy input from quantum pressure is masked by the more dominant compressible–to–incompressible transfer. These effects stem from the increased system size and partial spin-down, both of which hinder vortex expulsion and prolong the turbulent dynamics.
In the present model, certain aspects, such as crust potentials and related parameters, are treated phenomenologically, reflecting the diversity of existing models. Although a universal scaling law for glitch behavior across condensate sizes remains elusive, our simulations reveal qualitative consistency with turbulent scaling in larger condensates subjected to partial spin-down. Under these conditions, however, the contribution of quantum pressure to incompressible energy injection is notably reduced. This observation suggests that increased vortex expulsion enhances quantum pressure-driven energy transfer, highlighting the interplay between vortex dynamics and energy injection mechanisms.